3. Find all eigenvalues and eigenvectors of the matrix -2 0 -1 0 2 0 2 1 -2
Answer #1

A=\begin{bmatrix} -2&0 &2 \\ -1&0 & 1\\ 2&0 & -2 \end{bmatrix}

|A-\lambda I|=0

-\lambda ^3-4\lambda ^2=0

Eigen values

\lambda =0,-4

Eigen vector

For \lambda =0

\begin{bmatrix} -2& 0& 2\\ -1&0 &1 \\ 2 & 0 & -2 \end{bmatrix}

R1=R1/-2

\begin{bmatrix} 1& 0& -1\\ -1&0 &1 \\ 2 & 0 & -2 \end{bmatrix}

R2+R1=R2

\begin{bmatrix} 1& 0& -1\\ 0&0 &0 \\ 2 & 0 & -2 \end{bmatrix}

R3-2R1=R3

\begin{bmatrix} 1& 0& -1\\ 0&0 &0 \\ 0 & 0 & 0 \end{bmatrix}

x1=x3

x2=x2

x3=x3

V1=\begin{bmatrix} 0\\ 1\\ 0 \end{bmatrix}

V2=\begin{bmatrix} 1\\ 0\\ 1\end{bmatrix}

For \lambda =-4 ,A=\begin{bmatrix} 2& 0& 2\\ -1&4 &1 \\ 2 & 0 & 2 \end{bmatrix}

R1/2=R1

R3/2=R3

=\begin{bmatrix} 1& 0& 1\\ -1&4 &1 \\ 1 & 0 & 1 \end{bmatrix}

R3-R1=R3

R2=R2+R1

=\begin{bmatrix} 1& 0& 1\\ 0&4 &2\\ 0 & 0 & 0 \end{bmatrix}

R2=R2/4

=\begin{bmatrix} 1& 0& 1\\ 0&1&1/2\\ 0 & 0 & 0 \end{bmatrix}

x1+x3=0

x1=-x3

x2 +(x3)/2=0

x2=-(x3)/2

x3=x3

V3=\begin{bmatrix} -1\\ -1/2\\ 1 \end{bmatrix}